# Optimal Dispatch of Electricity and Water in Renewable-Integrated Desalination Plants

### v1

@@ Abstract
We develop a mathematical framework for the optimal dispatch of flexible water desalination plants (WDPs) as hybrid generator-load resources. WDPs integrate thermal generation, membrane-based controllable loads, and renewable energy sources, offering unique operational flexibility for power system operations. They can simultaneously participate in two markets: selling desalinated water to a water utility, and bidirectionally transacting electricity with the grid based on their net electricity demand. We formulate the dispatch decision problem of a profit-maximizing WDP, capturing operational, technological, and market-based coupling between water and electricity flows. The threshold-based structure we derive provides computationally tractable coordination suitable for large-scale deployment, offering operational insights into how thermal generation and membrane-based loads complementarily provide continuous bidirectional flexibility. The thresholds are analytically characterized in closed form as explicit functions of technology and tariff parameters. We examine how small changes in the exogenous tariff and technology parameters affect the WDP’s profit. Extensive simulations illustrate the optimal WDP’s operation, profit, and water-electricity exchange, demonstrating significant improvements relative to benchmark algorithms.
@@ I Introduction
The ongoing transformation of power systems, driven by increasing penetration of variable renewables, creates growing demand for flexible resources that can provide bidirectional grid services. Water desalination plants (WDPs), [^1] By WDP, we also refer to water treatment plants. The two differ only in terms of the salinity coefficients considered [Oikonomou&Parvania:20TSG]. representing approximately 5-12% of total electricity consumption in Middle East countries [Siddiqi&Anadon:11AE] with 7% per annum global capacity growth since 2010 [Eke&Yusuf&Giwa&Sodiq:20Desalination], are evolving from passive to active market participants capable of dynamically adjusting operations in response to price signals and renewable energy availability. Unlike conventional generators that can only supply power or loads that can only reduce consumption, modern WDPs integrating both thermal and renewable generation and membrane-based loads function as hybrid generator-load resources with continuous bidirectional flexibility.
With global freshwater demand projected to exceed available supply by 40% by 2030 [TurningTheTide:23GlobalWaterCommission], WDPs have become an increasingly vital component of sustainable water infrastructure in various parts of the world. As the demand for seawater and wastewater treatment continues to grow, so does the need for efficient and sustainable operation of WDPs. These facilities are inherently energy-intensive, consuming substantial amounts of energy to drive both thermal- and membrane-based desalination processes [Elimelech:2011Science]. Consequently, there is growing interest in integrating desalination operations with renewable energy sources and advanced dispatching strategies to reduce operational costs and environmental impact [Ghaffour:2015Desalination].
In parallel, the ongoing transformation of power systems, driven by increasing penetration of renewables and the rise of decentralized energy resources, has given rise to novel opportunities and challenges at the water-energy nexus. In this context, WDPs are evolving from passive energy consumers to active participants in electricity markets, capable of dynamically adjusting their operations in response to electricity price signals and renewable energy availability. Co-optimizing water and power dispatching, a global research priority in the water-energy nexus [ONeiletal:24OAJPE], is crucial for ensuring economic viability and grid-supportive behavior.
This paper develops an analytical framework for the optimal dispatch of WDPs operating as hybrid generator-load resources (Fig.1). A key challenge lies in utilizing the complementarities between the load-generator resources within the plant, in addition to optimally dispatching both water and electricity transactions with water and electric utilities.
@@ I-A Related Work
One of the earliest examinations of water-power systems is Gleick’s seminal review [Gleick:94ARRR], which highlighted the fundamental interdependence between water and power systems: energy is required for water extraction, treatment, and transport, while water is essential for electricity generation and fuel production. This foundational perspective has since motivated a broad literature on modeling, coordinating, and optimizing coupled water-energy infrastructures.
@@
In energy-abundant regions, TDPs often operate alongside power generation, producing freshwater as the primary output and electricity as a by-product. These colocated systems leverage shared thermal and electrical infrastructure, resulting in strongly coupled operational behavior. For instance, [Santhosh&Farid&Toumi:14AE] propose dynamic, physics-based models of such integrated systems, emphasizing the value of coordinated control between thermal power and thermal desalination units.
@@
Reverse osmosis (RO) has emerged as the dominant desalination technology, favored for its modularity and superior energy efficiency. Recent estimates suggest that RO accounts for approximately 69% of global installed desalination capacity [Eke2020]. As RO units rely predominantly on electricity from the grid, they can function as substantial and flexible loads within the power system. Several recent studies [Ghaithan2022, Qudah2024, Almehizia&Almasri&Hussein&Ehsani:19TSG, Guoetal:16TSG] explore the integration of RODPs with variable renewable energy sources (e.g., solar and wind), primarily through simulation-based energy management frameworks designed to reduce operational costs, emissions, and improve grid compatibility. In particular, [Guoetal:16TSG] present an energy management system for a standalone wind-powered desalination microgrid, highlighting coordinated control strategies for matching variable generation to water demand. However, these contributions remain largely simulation-based and typically treat the power system as an exogenous input, without explicitly modeling the endogenous coupling between the water and energy sectors. More recent efforts [MoazeniAPEN2020, Moazeni&Khazaei&Asrari:21TSG, AlAwamiTPS2022] address this gap by developing numerical co-optimization models for water-energy microgrids and joint RO-power dispatching under dynamic demand, grid constraints. These studies also explore the role of RO in demand response and grid services. The work in [Mohammadi&Ardakani&Abdullah&Heydt&Thomas:19TPS] formulates a security-constrained unit commitment problem to coordinate the operation of RO-only desalination plants and power plants and showed a reduction in the system’s operating costs by 5% to 8% under the IEEE 118-bus system.
@@
Despite these advances, existing approaches are predominantly computational, with limited analytical insight into structural properties of the optimal dispatching problem, an issue that our work aims to address.
@@
Beyond RO-power coordination, recent work expands to decentralized water-energy nexus markets and flexible cross-sector operations. [ZhaoetalBlockchain:23TSG] propose a blockchain-based transactive framework for interconnected water-energy hubs, showing performance gains under spatial-temporal uncertainty. Similarly, [CaoetalWatershed:23TSG] develops a watershed-electricity model that exploits hydrological flexibility, through dynamic river conditions and pump dispatching, to support distribution network operations. While demonstrating the benefits of decentralized coordination, these studies do not consider colocating and the unique operating regimes of joint RO and TDP systems.
@@
The studies in [Oikonomou&Parvania:20TSG, Zamzam&Emiliano&Zhao&Taylor&Sidiropoulos:19TCNS, Gu&Sioshansi:25PESGM] propose models for coordinating water and power distribution systems to minimize operational costs of meeting water and electricity demands by leveraging variable speed pumps and water tanks. However, these models do not incorporate colocated renewable energy sources or bidirectional energy transactions between WDPs and the electric grid, limiting their ability to fully optimize integrated resource management.
@@
Although significant progress has been made, most existing work models RODP and TDP separately, without capturing their distinct operational features or interactions with water and power systems, particularly when colocated with renewables. This siloed treatment overlooks opportunities for coordinated real-time dispatching and efficient resource utilization. In addition, existing studies are predominantly numerical- and simulation-based, whereas we develop an analytical framework that enables scalable solutions with structural insights, including operational intuitions and comparative statics.
@@ I-B Main Contributions
Our main contribution is the analytical characterization of optimal electricity-water dispatch for WDPs operating as hybrid generator-load resources, enabling efficient grid integration while maintaining water production requirements. The threshold-based structure we derive provides computationally tractable coordination suitable for large-scale deployment, offering operational insights into how thermal generation and membrane-based loads complementarily provide continuous bidirectional flexibility.
@@
The most pertinent scenario arises when the marginal value of electricity when water is produced via RO exceeds the power export price but remains below the import price, i.e., $f^{\prime}_{r}\pi^{w}\in[\pi^{-},\pi^{+}]$. Under this condition, we demonstrate that the optimal dispatch for TDP and RODP exhibits four-threshold and two-threshold structures, respectively. As onsite renewables increase, both TDP and RODP dynamically adjust to absorb surplus generation, expanding the power-balanced (NZ) mode while shrinking the power-importing (IM) and exporting (EX) modes.
In addition to the general structure of the solution, we analyze the optimal WDP dispatch in cases of single desalination technology. Also, we show the effect of the exogenous water and electricity tariff and technology parameters on the WDP’s maximum profit. We provide extended simulation results illustrating the optimal WDP’s profit, dispatch, and patterns of water and electricity exchange compared to benchmarks.
Figure 2: Optimal WDP dispatch across electricity operating modes (IM: import, NZ: net-zero, EX: export).
@@ I-C Major Mathematical Symbols and Paper Organization
$\Gamma_{\sf{IM}},\Gamma_{\sf{EX}}$ Mode-defining optimal dispatch thresholds.
$\Gamma_{\sf{NZ_{1}}},\Gamma_{\sf{NZ_{2}}}$ Internal optimal dispatch thresholds.
@@ II WDP Framework and Dispatch Problem
From a grid operator perspective, the WDP represents a controllable hybrid resource capable of both generation and consumption. Unlike conventional resources with unidirectional power flow, the WDP can smoothly transition between grid import, balanced, and export states through coordinated control of TDP and RODP units. This section formalizes the WDP’s operational characteristics and dispatch problem.
We consider the problem of a WDP that jointly optimizes water and electricity dispatch to maximize its profits from transacting the two commodities with water and electric utilities (Fig.1). The WDP integrates thermal and membrane-based desalination units, colocated with renewables on-site. It engages in a unilateral transaction with the water utility, whereby it supplies water at an exogenously specified price, which may vary over time. In contrast, its interaction with the electricity utility is bilateral: the WDP purchases electric power when its demand exceeds local generation and exports surplus power otherwise. This arrangement is structurally analogous to virtual power plants [Koraki&Strunz:18TPS] and grid-connected microgrids [Jiang&Xue&Geng:13TPS]. In both the water and electricity markets, the WDP is modeled as a price-taking agent, with its dispatch affecting quantities but not market-clearing prices.
@@
The WDP’s operational timescale is set by its measurement and dispatching frequency, while transactions with water and electric utilities follow their billing periods (e.g., 15 min or hourly). To streamline the exposition, we assume the WDP’s dispatch period matches the utilities’ billing period, allowing a single time-step formulation.
Next, we present the modeling of the TDP, RODP, payment functions to the water and electricity utilities, and the WDP dispatch problem.
@@ II-A1 TDP
The TDP, therefore, consumes fuel, denoted by $p_{h}$, and produces both water and electricity, which we denote by $w_{h}\in\mathbb{R}_{+}$ and $q_{h}\in\mathbb{R}_{+}$, respectively. We adopt a standard linear function of TDP desalinated water $w_{h}=f_{h}(p_{h})=\alpha_{h}p_{h}$, where $\alpha_{h}\in\mathbb{R}_{+}$ is the thermal conversion factor (m3/BTU) that represents the consumed fuel (BTU) for each cubic meter of desalinated water (m3) [Alnory&Beltagy:RE14]. We use $f_{h}^{\prime}$ to denote the derivative of $f_{h}$ with respect to $p_{h}$, i.e., $f_{h}^{\prime}:=\frac{\partial f_{h}(p_{h})}{\partial p_{h}}$.
The TDP generates electricity as a by-product of water desalination according to the relation $w_{h}=\eta_{h}q_{h}$, where $\eta_{h}\in\mathbb{R}_{++}$ is the water-to-electricity production ratio (m3/kW) [Santhosh&Farid&Toumi:14AE]. Define $\beta_{h}:=\alpha_{h}/\eta_{h}\in\mathbb{R}_{+}$, which is the thermal conversion factor (kW/BTU) from fuel to electricity.
The TDP water flowrate is bounded by the maximum $\overline{w}_{h}\in\mathbb{R}_{+}$ and minimum $\underline{w}_{h}\in\mathbb{R}_{+}$ operational limits that ensure stable operation, i.e., $w_{h}\in[\underline{w}_{h},\overline{w}_{h}]$.
@@ II-A2 RODP
The RODP consumes electricity to force pressure seawater (or wastewater) molecules through a semi-permeable membrane under high pressure, effectively filtering out salt and impurities [Mulder:96Springer].
@@
The water production of the RODP is bounded from above by the maximum $\overline{w}_{r}\in\mathbb{R}_{+}$ flowrate and from below by the minimum $\underline{w}_{r}\in\mathbb{R}_{+}$ operational limits, i.e., $w_{r}\in[\underline{w}_{r},\overline{w}_{r}]$.
@@ II-A3 Renewable Energy Sources
The plant has local renewables whose aggregate generation is denoted by $g\in\mathbb{R}_{+}$. We assume perfect forecasts of renewable generation. This is a simplifying assumption supported by the smoothing effect of coupling renewables with energy storage and the increasing accuracy of machine learning–based forecasting methods [Hong&Pinson:16book, Chen&Wang&Kirschen&Zhang:18TPS]. Capital costs of renewable installations are excluded from the analysis, as we focus on short-run operational dispatching rather than long-term capacity planning.
@@ II-B2 Electricity Payment
where the parameters $(\pi^{+},\pi^{-})\in\mathbb{R}_{+}$ represent the import and export prices, respectively, and $\pi^{0}\in\mathbb{R}$ the non-volumetric fixed charge. To avoid utility death spirals and risk-free arbitrage by the WDP, we assume $\pi^{+}\geq\pi^{-}$. A special case of (3) is when the WDP faces a single price ($\pi^{+}=\pi^{-}$), e.g., an WDP facing LMPs as in competitive electricity markets.
From (3), the WDP is a power-importer (IM) if $z>0$, power-exporter (EX) if $z<0$, and power-balanced (NZ) if $z=0$.
Under a multi-interval formulation, water and electricity payments naturally generalize to time-varying prices, such as time-of-use or real-time tariffs, which may be updated monthly or seasonally [Alahmed&Tong:22IEEETSG]. Alternatively, utilities can establish long-term contracts with desalination plants that fix import and export prices.

### v2

# Optimal Scheduling of Electricity and Water in Renewable-Colocated Desalination Plants

@@ Abstract
We develop a mathematical framework for the optimal scheduling of flexible water desalination plants (WDPs) as hybrid generator-load resources. WDPs integrate thermal generation, membrane-based controllable loads, and renewable energy sources, offering unique operational flexibility for power system operations. They can simultaneously participate in two markets: selling desalinated water to a water utility, and bidirectionally transacting electricity with the grid based on their net electricity demand. We formulate the scheduling decision problem of a profit-maximizing WDP, capturing operational, technological, and market-based coupling between water and electricity flows. The threshold-based structure we derive provides computationally tractable coordination suitable for large-scale deployment, offering operational insights into how thermal generation and membrane-based loads complementarily provide continuous bidirectional flexibility. The thresholds are analytically characterized in closed form as explicit functions of technology and tariff parameters. We examine how small changes in the exogenous tariff and technology parameters affect the WDP’s profit. Extensive simulations illustrate the optimal WDP’s operation, profit, and water-electricity exchange, demonstrating significant improvements relative to benchmark algorithms.
@@ I Introduction
The ongoing transformation of power systems, driven by increasing penetration of variable renewables, creates growing demand for flexible resources that can provide bidirectional grid services. Water desalination plants (WDPs), [^1] By WDP, we also refer to water treatment plants. The two differ only in terms of the salinity coefficients considered [24]. representing approximately 5-12% of total electricity consumption in Middle East countries [27] with 7% per annum global capacity growth since 2010 [8], are evolving from passive to active market participants capable of dynamically adjusting operations in response to price signals and renewable energy availability. Unlike conventional generators that can only supply power or loads that can only reduce consumption, modern WDPs integrating both thermal and renewable generation and membrane-based loads function as hybrid generator-load resources with continuous bidirectional flexibility.
With global freshwater demand projected to exceed available supply by 40% by 2030 [14], WDPs have become an increasingly vital component of sustainable water infrastructure in various parts of the world. As the demand for seawater and wastewater treatment continues to grow, so does the need for efficient and sustainable operation of WDPs. These facilities are inherently energy-intensive, consuming substantial amounts of energy to drive both thermal- and membrane-based desalination processes [9]. Consequently, there is growing interest in integrating desalination operations with renewable energy sources and advanced scheduling strategies to reduce operational costs and environmental impact [11].
In parallel, the ongoing transformation of power systems, driven by increasing penetration of renewables and the rise of decentralized energy resources, has given rise to novel opportunities and challenges at the water-energy nexus. In this context, WDPs are evolving from passive energy consumers to active participants in electricity markets, capable of dynamically adjusting their operations in response to electricity price signals and renewable energy availability. Co-optimizing water and power scheduling, a global research priority in the water-energy nexus [23], is crucial for ensuring economic viability and grid-supportive behavior.
This paper develops an analytical framework for the optimal scheduling of WDPs operating as hybrid generator-load resources (Fig.1). A key challenge lies in utilizing the complementarities between the load-generator resources within the plant, in addition to optimally scheduling both water and electricity transactions with water and electric utilities.
@@ I-A Related Work
One of the earliest examinations of water-power systems is Gleick’s seminal review [13], which highlighted the fundamental interdependence between water and power systems: energy is required for water extraction, treatment, and transport, while water is essential for electricity generation and fuel production. This foundational perspective has since motivated a broad literature on modeling, coordinating, and optimizing coupled water-energy infrastructures.
@@
In energy-abundant regions, TDPs often operate alongside power generation, producing freshwater as the primary output and electricity as a by-product. These colocated systems leverage shared thermal and electrical infrastructure, resulting in strongly coupled operational behavior. For instance, [26] propose dynamic, physics-based models of such integrated systems, emphasizing the value of coordinated control between thermal power and thermal desalination units.
@@
Reverse osmosis (RO) has emerged as the dominant desalination technology, favored for its modularity and superior energy efficiency. Recent estimates suggest that RO accounts for approximately 69% of global installed desalination capacity [8]. As RO units rely predominantly on electricity from the grid, they can function as substantial and flexible loads within the power system. Several recent studies [12, 25, 4, 16] explore the integration of RODPs with variable renewable energy sources (e.g., solar and wind), primarily through simulation-based energy management frameworks designed to reduce operational costs, emissions, and improve grid compatibility. In particular, [16] present an energy management system for a standalone wind-powered desalination microgrid, highlighting coordinated control strategies for matching variable generation to water demand. However, these contributions remain largely simulation-based and typically treat the power system as an exogenous input, without explicitly modeling the endogenous coupling between the water and energy sectors. More recent efforts [19, 20, 10] address this gap by developing numerical co-optimization models for water-energy microgrids and joint RO-power dispatching under dynamic demand, grid constraints. These studies also explore the role of RO in demand response and grid services. The work in [21] formulates a security-constrained unit commitment problem to coordinate the operation of RO-only desalination plants and power plants and showed a reduction in the system’s operating costs by 5% to 8% under the IEEE 118-bus system.
@@
Despite these advances, existing approaches are predominantly computational, with limited analytical insight into structural properties of the optimal scheduling problem, an issue that our work aims to address.
@@
Beyond RO-power coordination, recent work expands to decentralized water-energy nexus markets and flexible cross-sector operations. [30] propose a blockchain-based transactive framework for interconnected water-energy hubs, showing performance gains under spatial-temporal uncertainty. Similarly, [6] develops a watershed-electricity model that exploits hydrological flexibility, through dynamic river conditions and pump scheduling, to support distribution network operations. While demonstrating the benefits of decentralized coordination, these studies do not consider colocating and the unique operating regimes of joint RO and TDP systems.
@@
The studies in [24, 29, 15] propose models for coordinating water and power distribution systems to minimize operational costs of meeting water and electricity demands by leveraging variable speed pumps and water tanks. However, these models do not incorporate colocated renewable energy sources or bidirectional energy transactions between WDPs and the electric grid, limiting their ability to fully optimize integrated resource management.
@@
Although significant progress has been made, most existing work models RODP and TDP separately, without capturing their distinct operational features or interactions with water and power systems, particularly when colocated with renewables. This siloed treatment overlooks opportunities for coordinated real-time scheduling and efficient resource utilization. In addition, existing studies are predominantly numerical- and simulation-based, whereas we develop an analytical framework that enables scalable solutions with structural insights, including operational intuitions and comparative statics.
@@ I-B Main Contributions
Our main contribution is the analytical characterization of the optimal electricity-water scheduling for WDPs operating as hybrid generator-load resources, enabling efficient grid integration while maintaining water production requirements. The threshold-based structure we derive provides computationally tractable coordination suitable for large-scale deployment, offering operational insights into how thermal generation and membrane-based loads complementarily provide continuous bidirectional flexibility.
@@
The most pertinent scenario arises when the marginal value of electricity when water is produced via RO exceeds the power export price but remains below the import price, i.e., $f^{\prime}_{r}\pi^{w}\in[\pi^{-},\pi^{+}]$. Under this condition, we demonstrate that the optimal schedules for TDP and RODP exhibits four-threshold and two-threshold structures, respectively. As onsite renewables increase, both TDP and RODP dynamically adjust to absorb surplus generation, expanding the power-balanced (NZ) mode while shrinking the power-importing (IM) and exporting (EX) modes.
In addition to the general structure of the solution, we analyze the optimal WDP dispatch in cases of single desalination technology. Also, we show the effect of the exogenous water and electricity tariff and technology parameters on the WDP’s maximum profit. We provide extended simulation results illustrating the optimal WDP’s profit, scheduling, and patterns of water and electricity exchange compared to benchmarks.
Figure 2: Optimal WDP water dispatch across electricity operating modes (IM: import, NZ: net-zero, EX: export).
@@ I-C Major Mathematical Symbols and Paper Organization
$\Gamma_{\sf{IM}},\Gamma_{\sf{EX}}$ Mode-defining optimal scheduling thresholds.
$\Gamma_{\sf{NZ_{1}}},\Gamma_{\sf{NZ_{2}}}$ Internal optimal scheduling thresholds.
@@ II WDP Framework and Dispatch Problem
From a grid operator perspective, the WDP represents a controllable hybrid resource capable of both generation and consumption. Unlike conventional resources with unidirectional power flow, the WDP can smoothly transition between grid import, balanced, and export states through coordinated control of TDP and RODP units. This section formalizes the WDP’s operational characteristics and scheduling problem.
We consider the problem of a WDP that jointly optimizes water and electricity schedules to maximize its profits from transacting the two commodities with water and electric utilities (Fig.1). The WDP integrates thermal and membrane-based desalination units, colocated with renewables on-site. It engages in a unilateral transaction with the water utility, whereby it supplies water at an exogenously specified price, which may vary over time. In contrast, its interaction with the electricity utility is bilateral: the WDP purchases electric power when its demand exceeds local generation and exports surplus power otherwise. This arrangement is structurally analogous to virtual power plants [18] and grid-connected microgrids [17]. In both the water and electricity markets, the WDP is modeled as a price-taking agent, with its dispatch affecting quantities but not market-clearing prices.
@@
The WDP’s operational timescale is set by its measurement and scheduling frequency, while transactions with water and electric utilities follow their billing periods (e.g., 15 min or hourly). To streamline the exposition, we assume the WDP’s scheduling period matches the utilities’ billing period, allowing a single time-step formulation.
Next, we present the modeling of the TDP, RODP, payment functions to the water and electricity utilities, and the WDP scheduling problem.
@@ II-A1 TDP
The TDP, therefore, consumes fuel, denoted by $p_{h}$, and produces both water and electricity, which we denote by $w_{h}\in\mathbb{R}_{+}$ and $q_{h}\in\mathbb{R}_{+}$, respectively. We adopt a standard linear function of TDP desalinated water $w_{h}=f_{h}(p_{h})=\alpha_{h}p_{h}$, where $\alpha_{h}\in\mathbb{R}_{+}$ is the thermal conversion factor (m3/BTU) that represents the consumed fuel (BTU) for each cubic meter of desalinated water (m3) [1]. We use $f_{h}^{\prime}$ to denote the derivative of $f_{h}$ with respect to $p_{h}$, i.e., $f_{h}^{\prime}:=\frac{\partial f_{h}(p_{h})}{\partial p_{h}}$.
The TDP generates electricity as a by-product of water desalination according to the relation $w_{h}=\eta_{h}q_{h}$, where $\eta_{h}\in\mathbb{R}_{++}$ is the water-to-electricity production ratio (m3/kW) [26]. Define $\beta_{h}:=\alpha_{h}/\eta_{h}\in\mathbb{R}_{+}$, which is the thermal conversion factor (kW/BTU) from fuel to electricity.
The TDP water flowrate is bounded by the maximum $\overline{w}_{h}\in\mathbb{R}_{+}$ and minimum $\underline{w}_{h}\in\mathbb{R}_{+}$ operating limits that ensure stable operation, i.e., $w_{h}\in[\underline{w}_{h},\overline{w}_{h}]$.
@@ II-A2 RODP
The RODP consumes electricity to force pressure seawater (or wastewater) molecules through a semi-permeable membrane under high pressure, effectively filtering out salt and impurities [22].
@@
The water production of the RODP is bounded from above by the maximum $\overline{w}_{r}\in\mathbb{R}_{+}$ flowrate and from below by the minimum $\underline{w}_{r}\in\mathbb{R}_{+}$ operating limits, i.e., $w_{r}\in[\underline{w}_{r},\overline{w}_{r}]$.
@@ II-A3 Renewable Energy Sources
The plant has local renewables whose aggregate generation is denoted by $g\in\mathbb{R}_{+}$. We assume perfect forecasts of renewable generation. This is a simplifying assumption supported by the smoothing effect of coupling renewables with energy storage and the increasing accuracy of machine learning-based forecasting methods [28, 7]. Capital costs of renewable installations are excluded from the analysis, as we focus on short-run operational dispatching rather than long-term capacity planning.
@@ II-B2 Electricity Payment
where the parameters $(\pi^{+},\pi^{-})\in\mathbb{R}_{+}$ represent the import and export prices, respectively, and $\pi^{0}\in\mathbb{R}$ the non-volumetric fixed charge. To avoid utility death spirals and risk-free arbitrage by the WDP, we assume $\pi^{+}\geq\pi^{-}$. A special case of (3) is when the WDP faces a single price ($\pi^{+}=\pi^{-}$), e.g., an WDP facing LMPs as in competitive electricity markets. From (3), the WDP is a power-importer (IM) if $z>0$, power-exporter (EX) if $z<0$, and power-balanced (NZ) if $z=0$.
Under a multi-interval formulation, water and electricity payments naturally generalize to time-varying prices, such as time-of-use or real-time tariffs, which may be updated monthly or seasonally [3]. Alternatively, utilities can establish long-term contracts with desalination plants that fix import and export prices.